OpenStudy (hellomynameiskam):

What property is shown in the equation? 3 ∙ (6ab) = 18ab A. associative property of multiplication B. commutative property of multiplication C. identity property of multiplication D. reciprocal property of multiplication Need help b4 time runs out

OpenStudy (anonymous):

I would say associative :)

OpenStudy (anonymous):

We see here that there are two groups being multiplied together. One group is the 3 by itself, and the other is the 6ab. By associative property, we can say that the answer is the same no matter how the multiplicative terms are grouped. Therefore,\[3 \times (6ab)=(3 \times 6ab) =18ab\]

OpenStudy (hellomynameiskam):

Simplify. -(3/4) (-8y) (4/3) (6x) (-1) A. –48xy B. –16xy C. 16xy D. 32xy

OpenStudy (hellomynameiskam):

thanks

OpenStudy (anonymous):

This is simple multiplication, so I'm not actually going to do it for you. However, I will set it up so it looks a bit nicer \[(-\frac{ 3 }{ 4 })(\frac{ 4 }{ 3 })(-8y)(6x)(-1)=?\] We see that we have 3 negatives. Negative times negative is a positive. And a positive times another negative is negative. So since we know our final answer is going to be negative, we can focus solely on the numerical values and add the sign at the end. Now just multiply :)